Solve : (x^3+xy^2+a^2y) dx+(y^3+yx^2+a^2x) dy=0
1
Expert's answer
2020-11-22T16:43:14-0500

"\displaystyle
(x^3+xy^2+a^2y)\mathrm{d}x + (y^3+yx^2+a^2x)\mathrm{d}y=0\\


x^3\mathrm{d}x + xy^2\mathrm{d}x + a^2y\mathrm{d}x + y^3\mathrm{d}y + yx^2\mathrm{d}y + a^2x \mathrm{d}y =0\\


x^3 \mathrm{d}x + y^3 \mathrm{d}y + xy^2\mathrm{d}x + yx^2\mathrm{d}y + a^2(y\mathrm{d}x + x\mathrm{d}y)=0\\
x^3 \mathrm{d}x + y^3 \mathrm{d}y + xy(y\mathrm{d}x + x\mathrm{d}y) + a^2(y\mathrm{d}x + x\mathrm{d}y)=0\\
x^3 \mathrm{d}x + y^3 \mathrm{d}y + (xy + a^2)(y\mathrm{d}x + x\mathrm{d}y) =0\\


\textsf{Integrating both sides}\\

\int x^3\, \mathrm{d}x + \int y^3 \,\mathrm{d}y + \int\,(xy + a^2)\mathrm{d}(xy) = C\\

\textsf{Substitute}\, u = xy\,\textsf{in the last integral}.\\

\int x^3\, \mathrm{d}x + \int y^3 \,\mathrm{d}y + \int\,(u + a^2)\mathrm{d}u = C\\

\frac{x^4}{4} + \frac{y^4}{4} + \frac{u^2}{2} + a^2u = C\\

\frac{x^4}{4} + \frac{y^4}{4} + \frac{(xy)^2}{2} + a^2xy = C\\

x^4 + y^4 + 2(xy)^2 + a^2xy = C\\

\therefore(x^2 + y^2)^2 + a^2xy = C \,\,\textsf{is a solution to the ODE}\\

\textsf{Where}\, C\, \textsf{is an arbitrary constant.}"